Three years ago, you made a deposit in a savings account with a 4% APR. The bank pays interest monthly with no compounding. Today you have $1680 in the bank. What was your original deposit?
Answer #1

P = $1,500.00

Equation:
P = A / (1 + rt)


Calculation:
First, converting R percent to r a decimal
r = R/100 = 4%/100 = 0.04 per year.

Putting time into years for simplicity,
36 months / 12 months/year = 3 years

Solving our equation:
P = 1680 / ( 1 + (0.04 × 3)) = 1500
P = $1,500.00